# Vector valued estimates for matrix weighted maximal operators and product $\mathrm{BMO}$

### v2

@@ 4.2 Two-matrix weighted bounds for paraproducts
$\lesssim_{n,p,d}[W_{x_{1}}]_{A_{p},\mathcal{D}_{2}}^{\max\left\{p^{\prime},\frac{p}{2}+\frac{1}{p-1}\right\}}\int_{\mathbb{R}^{m}}|W_{x_{1}}(x_{2})^{1/p}f(x_{1},x_{2})|^{p}\,\mathrm{dm}(x_{2})$
$\lesssim_{p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{\max\left\{p^{\prime},\frac{p}{2}+\frac{1}{p-1}\right\}}\int_{\mathbb{R}^{m}}|W(x_{1},x_{2})^{1/p}f(x_{1},x_{2})|^{p}\,\mathrm{dm}(x_{2}).$
@@
Let $d,p,U,V$ and $B$ be as above. If $\Pi$ is any of the above defined mixed biparameter paraproducts, then there holds
$\|\Pi\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(V)}\lesssim\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)},$
where all implied constants depend only on $n,m,d,p,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$
@@
Throughout the proof ,$\lesssim$,$\gtrsim$,$\sim$ , ≳ , ∼ mean that all implied inequality constants depend only on $n,m,d,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$
We adapt the factorization trick from the proof of [Holmes_Petermichl_Wick_2018, Proposition 6.1]. We treat as an example $\Gamma_{\bm{\mathcal{D}},B}^{(10)},$ the proof for the other mixed paraproducts being similar or following by duality. We have
$\left|(\Gamma_{\bm{\mathcal{D}},B}^{(10)}f,g)\right|$ $=\bigg{|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}\frac{1}{\sqrt{|R_{2}|}}\langle B_{R}^{\varepsilon}\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}},g_{R}^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}\rangle\bigg{|}$
@@
$\Phi_{R}^{\varepsilon}\coloneq\frac{1}{\sqrt{|R_{2}|}}\sum_{\delta_{2}\in\mathcal{E}^{2}\setminus\{\varepsilon_{2}\}}g^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}_{R}\overline{\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}}}^{T},\quad R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \varepsilon\in\mathcal{E}.$
@@
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\lesssim\|f\|_{L^{p}(U)}\|g\|_{L^{p^{\prime}}(V^{\prime})}.$ (4.6)
We have
$\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\bigg{|}V(x)^{-1/p}\left(\frac{1}{\sqrt{|R_{2}|}}\sum_{\delta_{2}\in\mathcal{E}^{2}\setminus\{\varepsilon_{2}\}}g^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}_{R}\overline{\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}}}^{T}\right)\mathcal{U}_{R}\bigg{|}^{2}\frac{\mathbf{1}_{R}(x)}{|R|}$
$\leq 2^{m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}|V(x)^{-1/p}g^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}_{R}\overline{\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}}}^{T}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R_{1}|\cdot|R_{2}|^{2}}$
$\leq 2^{m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}|V(x)^{-1/p}g^{\varepsilon}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\cdot|\mathcal{U}_{R}\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}}|^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}$
@@
where
$S_{\bm{\mathcal{D}},V^{\prime}}g(x)\coloneq\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|V(x)^{-1/p}g_{R}^{\varepsilon}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{1/2}$
and
$[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},U}f(x)\coloneq\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}(\sup_{R_{1}\in\mathcal{D}^{1}}|\mathcal{U}_{R}\langle f^{\varepsilon_{2},2}_{R_{2}}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1}))^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{1/2},$
for all $x=(x_{1},x_{2})\in\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Therefore
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\leq\|S_{\bm{\mathcal{D}},V^{\prime}}g\|_{L^{p^{\prime}}(\mathbb{R}^{n+m})}\|[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},U}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}.$
It is proved in [DKPS2024, Lemmas 5.2] that
$\|S_{\bm{\mathcal{D}},V^{\prime}}g\|_{L^{p^{\prime}}(\mathbb{R}^{n+m})}\lesssim\|g\|_{L^{p^{\prime}}(V^{\prime})}.$
@@ 4.3 Two matrix weighted upper bounds for bicommutators
Two matrix weighted upper bounds for biparameter paraproducts lead naturally to two matrix weighed upper bounds for bicommutators. Here we overview very briefly the simplest case: that of bicommutators with Haar multipliers on $\mathbb{R}.$ The argument is a straightforward adaptation of the one in the (unweighted) $\mathrm{L}^{2}$ scalar valued case from [Blasco_Pott_2005].
@@
Let $B:\mathbb{R}^{2}\to\mathrm{M}_{d}(\mathbb{C})$ be a locally integrable function. For scalar valued locally integrable functions $b$ on $\mathbb{R}^{2}$ it is shown in [Blasco_Pott_2005] that
$[T^{1}_{\sigma_{1}},[T^{2}_{\sigma_{2}},b]]=[T^{1}_{\sigma_{1}},[T^{2}_{\sigma_{2}},\Lambda_{b}]],$ (4.7)
@@
$[T^{1}_{\sigma_{1}},[T^{2}_{\sigma_{2}},B]]=[T^{1}_{\sigma_{1}},[T^{2}_{\sigma_{2}},\Lambda_{B}]],$
@@
Let now $1<p<\infty$ and let $U,V$ be biparameter $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic $A_{p}$ weights on $\mathbb{R}\times\mathbb{R}.$ In the following estimates, all implied constants depend only on $d,p,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$ Using (2.6.3) and the well-known two matrix weighted bounds in the one parameter setting we deduce
@@ Appendix A Appendix
Through the article we have used the results of Bownik and Cruz-Uribe [Cruz_Uribe_Bownik_Extrapolation] in the complex setting. The results in that paper are stated in the real setting and it is not immediate that they work in the complex setting. Some modifications are necessary and some steps require proper justification. For instance, given a family of norms measurably parametrized (over $\mathbb{C}^{d}$), it is not obvious that one can assign them a reducing matrix and a complex John ellipsoid in a measurable way. This is shown in detail in [DKPS2024, Appendix A]. We devote this appendix to briefly discussing the necessary modifications and where to find the details to get the results of [Cruz_Uribe_Bownik_Extrapolation] for complex valued matrix weights.
@@ A.1 Convex sets and seminorms in the complex setting
We begin with the required changes in [Cruz_Uribe_Bownik_Extrapolation, Section 2]. In both the present work and that of Bownik and Cruz-Uribe, one considers symmetric sets. The difference is that we substitute real symmetric sets $E\subseteq\mathbb{R}^{d},$ that is sets $E$ such that $-u\in E$ for every $u\in E,$ by complex symmetric sets $E\subseteq\mathbb{C}^{d},$ i.e. sets $E$ for which $\lambda u\in E$ for every $u\in E$ and for any $\lambda\in\mathbb{C}$ with $|\lambda|=1.$ Let us note that the “symmetric convex sets” as we have defined them are precisely the balanced convex sets, meaning that $\lambda u\in E$, for all $u\in E$ and $\lambda\in\mathbb{C}$ with $|\lambda|\leq 1.$ Besides this, the rest of the definitions concerning convex bodies, including that of the Minkowski addition of sets, is the same. Of course, the definitions for seminorms are also the same, only that we consider them to be defined on the vector space $\mathbb{C}^{d}$ and to be complex homogeneous functions, which means that if $\rho$ is a seminorm on $\mathbb{C}^{d},$ then $\rho(\lambda v)=|\lambda|\rho(v)$ for any $v\in\mathbb{C}^{d}$ and $\lambda\in\mathbb{C}.$ Also, the fact that seminorms are in a one-to-one correspondence with absorbing convex (complex) symmetric bodies (the analogous result to [Cruz_Uribe_Extrapolation, Theorem 2.4]) also holds with the same proof (see [Rudin_Functional_Analysis_1991, Theorems 1.34 and 1.35]).
Dual seminorms, their properties and their relation to the polar of convex (complex) symmetric sets follow the same exposition in the vector space $\mathbb{C}^{d}$ as in [Cruz_Uribe_Bownik_Extrapolation]. Here one only needs to keep in mind that we substitute the real Euclidean product by the complex Hermitian product of $\mathbb{C}^{d}.$ For this reason, in this context the support function $h_{K}$ of a set $K\in\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ has to be defined as
@@
(cf. [Cruz_Uribe_Bownik_Extrapolation, Definition 2.12]). With this definition, the proof of [Cruz_Uribe_Bownik_Extrapolation, Theorem 2.11], which covers the properties of seminorms and their relation to convex bodies, needs to be adapted. For instance, to show that
$p_{(K_{1}+K_{2})^{\circ}}=p_{K_{1}^{\circ}}+p_{K_{2}^{\circ}},$
one first shows that
$p_{(K_{1}+K_{2})^{\circ}}\leq p_{K_{1}^{\circ}}+p_{K_{2}^{\circ}}$
by the triangle inequality. Then, the reverse inequality is proved by taking $\varepsilon>0$ arbitrary and $w_{1}\in K_{1}$ and $w_{2}\in K_{2}$ with $\langle v,w_{j}\rangle\geq h_{K_{j}}(v)-\varepsilon/2$ for $j=1,2$ (we can omit the moduli here because of complex symmetry). This argument has to be repeated through the rest of the statements of [Cruz_Uribe_Bownik_Extrapolation, Theorem 2.11]. The properties of weighted geometric means of norms follow the same explanation for norms defined on $\mathbb{C}^{d}.$
For the rest of the section, the facts about positive definite matrices follow by the same arguments, since the facts that the authors of [Cruz_Uribe_Bownik_Extrapolation] use hold in the complex setting (see [Bhatia_Positive_definite_matrices_2007, Chapter 6]). Here, one only needs to take into account that we consider hermitian matrices instead of symmetric ones and that we use unitary matrices instead of orthogonal ones.
@@ A.2 Convex-set valued functions in the complex setting
In this subsection we explain the necessary adaptations that have to be performed in Section 3 of [Cruz_Uribe_Bownik_Extrapolation] in the complex setting. In some places $\mathbb{R}^{d}$ needs to be replaced by $\mathbb{R}^{2d}$, which as a topological space, as a metric space, as a measure space and as a real vector space is the same as $\mathbb{C}^{d}$. In some other places, $\mathbb{R}^{d}$ needs to be replaced directly by $\mathbb{C}^{d}$ and “complex” versions of the ingredients of statements or proofs are necessary. We lay out the details below.
First of all, measurable maps $F:\Omega\to\mathcal{K}(\mathbb{C}^{d})$ from a positive, $\sigma$-finite, complete measure space $\Omega$ into the set $\mathcal{K}(\mathbb{C}^{d})$ of closed subsets of $\mathbb{C}^{d}$ are the same as measurable maps $F:\Omega\to\mathcal{K}(\mathbb{R}^{2d})$. The various characterizations of measurability for maps $F:\Omega\to\mathcal{K}(\mathbb{R}^{d})$ given in [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.2] with $\mathbb{R}^{d}$ as the underlying space continue to hold with $\mathbb{R}^{d}$ replaced by $\mathbb{R}^{2d}$, and thus $\mathbb{C}^{d}$ as the underlying space. Let us observe in particular that a measurable selection function $f:\Omega\to\mathbb{C}^{d}$ of a map $F:\Omega\to\mathcal{K}(\mathbb{C}^{d})$ is the same as a measurable selection function $f:\Omega\to\mathbb{R}^{2d}$ of the map $F:\Omega\to\mathcal{K}(\mathbb{R}^{2d})$.
Similarly, the measurability of closed convex hulls of countable unions and the measurability of countable intersections of convex body valued functions [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.3] hold in $\mathbb{R}^{2d}$ and therefore also in $\mathbb{C}^{d}.$ A similar remark applies to [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.5], which states the equivalence between the definition of closed set valued functions and the measurability as a mapping with respect to the Hausdorff topology.
Referring to [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.4] proving the measurability of the polar of a measurable map, one replaces $\mathbb{R}^{d}$ directly by $\mathbb{C}^{d}$. This entails that complex symmetric convex sets are considered, and that the real Euclidean product on $\mathbb{R}^{d}$ needs to be replaced by the complex Hermitian product on $\mathbb{C}^{d}$. A similar change applies to [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.6]. This result applies to measurable set valued functions $F$ taking values on the set of linear subspaces of $\mathbb{C}^{d},$ and it states that if $P(x)$ is the orthogonal projection from $\mathbb{C}^{d}$ to $F(x),$ the mapping $P\colon\Omega\rightarrow\mathrm{M}_{{d}}({\mathbb{C}})$ is also measurable. The necessary change is to replace $\mathbb{R}^{d}$ by $\mathbb{C}^{d}$, so that orthogonal projections and the Gram–Schmidt orthonormalization process are considered with respect to the Hermitian product on $\mathbb{C}^{d}$. Real linear subspaces are replaced by complex linear subspaces. At the last step of the proof of [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.6], a countable dense subset of $\mathbb{C}^{d}$ is needed, one can take $\mathbb{Q}(i)^{d}$ for example.
@@
Next, one needs the existence and properties of John ellipsoids for bounded complex symmetric convex subsets of $\mathbb{C}^{d}$. These are thoroughly established in [DKPS2024, Appendix A] and the distinction between complex and real ellipsoids is made precise. In fact, it is implicit in [DKPS2024, Subsection A.4.7] that a real ellipsoid in $\mathbb{R}^{2d}$ is a complex ellipsoid in $\mathbb{C}^{d}$ if and only if it is invariant under the real orthogonal map $\widetilde{L}_{z}$ for all $z\in\mathbb{C}$ with $|z|=1$. For the reader’s convenience we lay out the details below.
We identify $\mathbb{R}^{2d}$ with $\mathbb{C}^{d}$ in the natural way, namely by the map $R:\mathbb{C}^{d}\to\mathbb{R}^{2d}$ given by

### v3

@@ 4.2 Two-matrix weighted bounds for paraproducts
$\lesssim_{n,p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{\max\left\{p^{\prime},\frac{p}{2}+\frac{1}{p-1}\right\}}\int_{\mathbb{R}^{m}}|W(x_{1},x_{2})^{1/p}f(x_{1},x_{2})|^{p}\,\mathrm{dm}(x_{2}).$
@@
Let $d,p,U,V$ and $B$ be as above. Then there holds
$\|\Gamma_{\bm{\mathcal{D}},B}^{(10)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(V)}\lesssim\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)},$
where all implied constants depend only on $n,m,d,p,[U]_{A_{p},\bm{\mathcal{D}}}$ and $V]_{A_{p},\bm{\mathcal{D}}}.$
@@
Throughout the proof $\lesssim,\gtrsim,\sim$ mean that all implied inequality constants depend only on $n,m,d,[U]_{A_{p},\bm{\mathcal{D}}}$ and $V]_{A_{p},\bm{\mathcal{D}}}.$
We adapt the factorization trick from the proof of [23, Proposition 6.1]. We have
$\left|(\Gamma_{\bm{\mathcal{D}},B}^{(10)}f,g)\right|$ $=\bigg|\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\penalty 10000\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}\frac{1}{\sqrt{|R_{2}|}}\langle B_{R}^{\varepsilon}\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}},g_{R}^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}\rangle\bigg|$
@@
$\Phi_{R}^{\varepsilon}\coloneq\frac{1}{\sqrt{|R_{2}|}}\sum_{\delta_{2}\in\mathcal{E}^{2}\setminus\{\varepsilon_{2}\}}g^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}_{R}\overline{\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}}}^{T},\quad R\in\bm{\mathcal{D}},\penalty 10000\ \varepsilon\in\mathcal{E}.$
@@
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\lesssim\|f\|_{\mathrm{L}^{p}(U)}\|g\|_{\mathrm{L}^{p^{\prime}}(V^{\prime})}.$ (4.17)
We can bound $\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}$ by
$2^{m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\penalty 10000\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}|V(x)^{-1/p}g^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}_{R}\overline{\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}}}^{T}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R_{1}|\cdot|R_{2}|^{2}}$
$\leq 2^{m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\penalty 10000\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}|V(x)^{-1/p}g^{\varepsilon}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\cdot|\mathcal{U}_{R}\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}}|^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}$
@@
where $S_{\bm{\mathcal{D}},V^{\prime}}g$ and $\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},U}f$ are defined as previously. Therefore
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\leq\|S_{\bm{\mathcal{D}},V^{\prime}}g\|_{\mathrm{L}^{p^{\prime}}(\mathbb{R}^{n+m})}\|[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},U}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}.$
It is proved in [14, Lemmas 5.2] that
$\|S_{\bm{\mathcal{D}},V^{\prime}}g\|_{\mathrm{L}^{p^{\prime}}(\mathbb{R}^{n+m})}\lesssim\|g\|_{\mathrm{L}^{p^{\prime}}(V^{\prime})}.$
@@ 4.3 Two matrix weighted upper bounds for bicommutators
Two matrix weighted upper bounds for biparameter paraproducts lead naturally to two matrix weighted upper bounds for bicommutators. Here we overview very briefly the simplest case: that of bicommutators with Haar multipliers on $\mathbb{R}.$ The argument is a straightforward adaptation of the one in the (unweighted) $\mathrm{L}^{2}$ scalar valued case from [3].
@@
Let $B:\mathbb{R}^{2}\to\mathrm{M}_{d}(\mathbb{C})$ be a locally integrable function. For scalar valued locally integrable functions $b$ on $\mathbb{R}^{2}$ it is shown in [3] that
$T^{1}_{\sigma_{1}},[T^{2}_{\sigma_{2}},b]]=[T^{1}_{\sigma_{1}},[T^{2}_{\sigma_{2}},\Lambda_{b}]],$ (4.18)
@@
$T^{1}_{\sigma_{1}},[T^{2}_{\sigma_{2}},B]]=[T^{1}_{\sigma_{1}},[T^{2}_{\sigma_{2}},\Lambda_{B}]],$
@@
Let now $1<p<\infty$ and let $U,V$ be biparameter $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic $A_{p}$ weights on $\mathbb{R}\times\mathbb{R}.$ In the following estimates, all implied constants depend only on $d,p,[U]_{A_{p},\bm{\mathcal{D}}}$ and $V]_{A_{p},\bm{\mathcal{D}}}.$ Using (2.4) and the well-known two matrix weighted bounds in the one parameter setting [13] we deduce
@@ Appendix A Appendix
Through the article we have used the results of Bownik and Cruz-Uribe [5] in the complex setting. The results in that paper are stated in the real setting and it is not immediate that they work in the complex setting. Some modifications are necessary and some steps require proper justification. For instance, given a family of norms measurably parametrized (over $\mathbb{C}^{d}$), it is not obvious that one can assign them a reducing matrix and a complex John ellipsoid in a measurable way. This is shown in detail in [14, Appendix A]. We devote this appendix to briefly discussing the necessary modifications and where to find the details to get the results of [5] for complex valued matrix weights.
@@ A.1 Convex sets and seminorms in the complex setting
We begin with the required changes in [5, Section 2]. In both the present work and that of Bownik and Cruz-Uribe, one considers symmetric sets. The difference is that we substitute real symmetric sets $E\subseteq\mathbb{R}^{d},$ that is sets $E$ such that $u\in E$ for every $u\in E,$ by complex symmetric sets $E\subseteq\mathbb{C}^{d},$ i.e. sets $E$ for which $\lambda u\in E$ for every $u\in E$ and for any $\lambda\in\mathbb{C}$ with $|\lambda|=1.$ Of course, the definitions for seminorms are also the same, only that we consider them to be defined on the vector space $\mathbb{C}^{d}$ and to be complex homogeneous functions, which means that if $\rho$ is a seminorm on $\mathbb{C}^{d},$ then $\rho(\lambda v)=|\lambda|\rho(v)$ for any $v\in\mathbb{C}^{d}$ and $\lambda\in\mathbb{C}.$
Dual seminorms, their properties and their relation to the polar of convex (complex) symmetric sets follow the same exposition in the vector space $\mathbb{C}^{d}$ as in [5]. Here one only needs to keep in mind that we substitute the real Euclidean product by the complex Hermitian product of $\mathbb{C}^{d}.$ For this reason, in this context the support function $h_{K}$ of a set $K\in\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ has to be defined as
@@
(cf. [5, Definition 2.12]).
For the rest of the section, the facts about positive definite matrices follow by the same arguments, since the facts that the authors of [5] use hold in the complex setting (see [2, Chapter 6]).
@@ A.2 Convex-set valued functions in the complex setting
Here we briefly explain the main relevant adaptations that have to be performed in Section 3 of [5] in the complex setting. Most of the results and arguments in that work can be ported to the complex setting in a straightforward way. The general principle that this relies on is the fact that one can identify $\mathbb{C}^{d}$ with $\mathbb{R}^{2d},$ being both equivalent as topological spaces, as metric spaces, as measure spaces and as real vector spaces. This, of course, requires implicitly other minor modifications, such as substituting the real Euclidean scalar product by the Hermitian scalar product.
The application of this general principle together with the work in [5] already yields the necessary facts about the measurability of $\mathcal{K}(\mathbb{C}^{d})$ valued maps and selection functions. Statements about measurability of closed convex hulls of countable unions and of countable intersections of convex body valued functions follow similarly. Moreover, a similar remark holds for the measurability of the polar map.
@@
The point that is not immediate to adapt, and which is one of the keystones in the development of the results in [5] is the existence and properties of John ellipsoids for bounded complex symmetric convex subsets of $\mathbb{C}^{d}$. These are thoroughly established in [14, Appendix A] and the distinction between complex and real ellipsoids is made precise. In fact, it is implicit in [14, Subsection A.4.7] that a real ellipsoid in $\mathbb{R}^{2d}$ is a complex ellipsoid in $\mathbb{C}^{d}$ if and only if it is invariant under a certain family of real orthogonal maps $\widetilde{L}_{z}.$
Denote by $\mathbf{\overline{B}}_{\mathbb{R}^{2d}}$ the closed unit ball in $\mathbb{R}^{2d}$ and by $\mathbf{\overline{B}}_{\mathbb{C}^{d}}$ the closed unit ball in $\mathbb{C}^{d}.$ Recall that a real ellipsoid in $\mathbb{R}^{2d}$ is a subset $E$ of $\mathbb{R}^{2d}$ of the form $E=A\mathbf{\overline{B}}_{\mathbb{R}^{2d}}$ for some real linear map $A:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$. It can be seen that, for a real ellipsoid $E,$ the linear map $A$ with $E=A\mathbf{\overline{B}}_{\mathbb{R}^{2d}}$ can be taken to be positive semidefinite and the choice of $A$ is unique under this condition. Similarly, a complex ellipsoid in $\mathbb{C}^{d}$ is a subset $E$ of $\mathbb{C}^{d}$ of the form $E=A\mathbf{\overline{B}}_{\mathbb{C}^{d}}$ for some complex linear map $A:\mathbb{C}^{d}\to\mathbb{C}^{d}$. It can also be seen that for each complex ellipsoid $E$ there is a unique complex positive semidefinite linear map $A$ with $E=A\mathbf{\overline{B}}_{\mathbb{C}^{d}}.$ A sketch of the real case can be found in [19, page 304]. For completeness we include a detailed proof in the complex case, adapting the ideas in [19, page 304].
Lemma 14.
Let $A,C\in M_{d}(\mathbb{C})$. Consider the complex ellipsoid $E:=A\mathbf{\overline{B}}_{\mathbb{C}^{d}}$. Then, we have $E=C\mathbf{\overline{B}}_{\mathbb{C}^{d}}$ if and only if $AA^{\ast}=CC^{\ast}$. In particular, $H:=|A^{\ast}|=(AA^{\ast})^{1/2}$ is the unique positive semidefinite matrix such that $E=H\mathbf{\overline{B}}_{\mathbb{C}^{d}}$.
Proof.
Since $AA^{\ast}$ and $CC^{\ast}$ are both self-adjoint, we have $AA^{\ast}=CC^{\ast}$ if and only if
$\langle AA^{\ast}x,x\rangle=\langle CC^{\ast}x,x\rangle,\quad\forall x\in\mathbb{C}^{d},$
that is
$|A^{\ast}x|=|C^{\ast}x|,\quad\forall x\in\mathbb{C}^{d}.$
Observe that for all $x\in\mathbb{C}^{d}$ it holds
$|A^{\ast}x|$ $=\sup\{|\langle A^{\ast}x,y\rangle|:\penalty 10000\ y\in\mathbf{\overline{B}}_{\mathbb{C}^{d}}\}=\sup\{|\langle x,Ay\rangle|:\penalty 10000\ y\in\mathbf{\overline{B}}_{\mathbb{C}^{d}}\}$
$=\sup\{|\langle x,z\rangle|:\penalty 10000\ z\in A\mathbf{\overline{B}}_{\mathbb{C}^{d}}\}$
and similarly
$|C^{\ast}x|=\sup\{|\langle x,z\rangle|:\penalty 10000\ z\in C\mathbf{\overline{B}}_{\mathbb{C}^{d}}\}.$
Therefore, if $A\mathbf{\overline{B}}_{\mathbb{C}^{d}}=C\mathbf{\overline{B}}_{\mathbb{C}^{d}}$, then it is clear that $|A^{\ast}x|=|C^{\ast}x|$, for all $x\in\mathbb{C}^{d}$ and thus $AA^{\ast}=CC^{\ast}$.
Conversely, assume that $AA^{\ast}=CC^{\ast}$. Then $|A^{\ast}|=(AA^{\ast})^{1/2}=(CC^{\ast})^{1/2}=|C^{\ast}|$. By the polar decomposition we can write $A^{\ast}=U|A^{\ast}|$ and $C^{\ast}=V|C^{\ast}|$ for some unitary matrices $U,V$, therefore
$A\mathbf{\overline{B}}_{\mathbb{C}^{d}}=|A^{\ast}|U^{\ast}\mathbf{\overline{B}}_{\mathbb{C}^{d}}=|A^{\ast}|\mathbf{\overline{B}}_{\mathbb{C}^{d}}=|C^{\ast}|\mathbf{\overline{B}}_{\mathbb{C}^{d}}=|C^{\ast}|V^{\ast}\mathbf{\overline{B}}_{\mathbb{C}^{d}}=C\mathbf{\overline{B}}_{\mathbb{C}^{d}}.$
Finally, the last claim follows immediately from the fact that $HH^{\ast}=AA^{\ast}$. ∎
Let us denote by $R:\mathbb{C}^{d}\to\mathbb{R}^{2d}$ the map given by